# Mesh2NeRF: Direct Mesh Supervision for Neural Radiance Field Representation and Generation

### v1

@@ 4.1 Single Scene Fitting
We evaluate our method on scenes from Amazon Berkeley Objects (ABO) [19], Poly Haven Models [1] and Sketchfab Scenes [2]. The ABO dataset contains twelve realistic 3D models of household objects. Each object is rendered at 512512 pixels from viewpoints sampled on a sphere, covering the object’s surface area (90 views as input and 72 for testing). We also show results on six hyper-real models from the Poly Haven library, which consists of quality content designed by artists. Each scene is rendered at 512512 with 90 input and 120 test views. We use the Entrée du château des Bois Francs scene from Sketchfab to evaluate large scene-scale data. The used Sketchfab scene is also rendered at 512512 pixels with 90 input views and 120 test views. For Mesh2NeRF, we use the textured mesh from both sets by setting a fixed point light. We evaluate our approach with three different encodings and parametric data structures for storing trainable feature embeddings, i.e., the network part of NeRFF [40], TensoRF [12], and Instant NGP [42] with our Mesh2NeRF direct supervision using the corresponding mesh. We denote these three methods by Mesh2NeRF NeRF, Mesh2NeRF TensoRF, and Mesh2NeRF NGP, respectively.
@@
We present qualitative results in Fig. 4, illustrating the superiority of our method over NeRF baselines in representing single scenes on challenging Coffee Cart from Poly Haven and Entrée du château des Bois Francs from Sketchfab. Mesh2NeRF TensoRF supasses the original TensoRF, indicating that Mesh2NeRF supervision contributes to the neural network’s ability to learn accurate scene representation. Mesh2NeRF NGP also outperforms Instant NGP, exhibiting fewer artifacts and providing more precise renderings. For additional implementation details, results and analysis, please refer to the supplemental material.
@@ 0.B.1 More Results on Single Scene Fitting
More qualitative comparsions. We present qualitative results from the Sketchfab dataset in Fig.10. Renderings from two views for each object are compared with NeRF baselines, accompanied by corresponding PSNR values. Similarly, Fig.11 showcases qualitative results on the ABO dataset, with renderings from two views for each object compared alongside corresponding LPIPS values. These figures demonstrate that Mesh2NeRF NeRF outperforms NeRF baselines.

### v2

@@ 4.1 Single Scene Fitting
We evaluate our method on scenes from Amazon Berkeley Objects (ABO) [20], Poly Haven Models [2], Sketchfab Scenes [3] and the Country-Kitchen scene [1]. The ABO dataset contains twelve realistic 3D models of household objects. Each object is rendered at 512512 pixels from viewpoints sampled on a sphere, covering the object’s surface area (90 views as input and 72 for testing). We also show results on six hyper-real models from the Poly Haven library, which consists of quality content designed by artists. Each scene is rendered at 512512 with 90 input and 120 test views. We use the Country-Kitchen scene to evaluate complex indoor scene under realistic lighting and material effect. It is also rendered at 512512 pixels with 1,208 input views and 205 test views. For Mesh2NeRF, we use the textured mesh from both sets by setting a fixed point light. We evaluate our approach with three different encodings and parametric data structures for storing trainable feature embeddings, i.e., the network part of NeRFF [41], TensoRF [13], and Instant NGP [43] with our Mesh2NeRF direct supervision using the corresponding mesh. We denote these three methods by Mesh2NeRF NeRF, Mesh2NeRF TensoRF, and Mesh2NeRF NGP, respectively.
@@
We present qualitative results in Fig. 4, illustrating the superiority of our method over NeRF baseline (Instant NGP) in representing single scenes on challenging Country-Kitchen. Mesh2NeRF NGP exhibiting fewer artifacts and providing more precise renderings. For additional implementation details, results and analysis, please refer to the supplemental material.
@@ 0.B.1 More Results on Single Scene Fitting
Lighting model. We apply Phong model in our experiments, while Mesh2NeRF can accommodate any BRDF and lighting. The only adjustment required is to change the function computing the color value of the ray hit point. For instance, if Spherical Harmonics lighting replaces the point lighting, the analytic solution (as illustrated in Fig. 3) can also achieve an SSIM over 0.99.
Baseline with geometry information. We compare NeRF baseline with and without GT depth supervision during fitting on the Sketchfab scene Entrée du château des Bois Francs. Results (Table 5) show improvements in Instant NGP with depth supervision in both dense (same as Sec. 4.1) and sparse (30 views) settings. More improvement from depth is gained in the limited views setup (+0.44 PSNR vs. +0.21). Even with depth, Instant NGP is still worse than Mesh2NeRF NGP.
| Method | Depth loss | PSNR |
| Instant NGP (limited views) | No | 15.75 |
| Yes | 16.19 (+0.44) | |
| Instant NGP (dense views) | No | 19.51 |
| Yes | 19.72 (+0.21) | |
| Mesh2NeRF NGP | - | 20.42 |
More qualitative comparsions. We present qualitative results in Fig. 10, illustrating the superiority of our method over NeRF baselines in representing single scenes on challenging Coffee Cart from Poly Haven and Entrée du château des Bois Francs from Sketchfab. Mesh2NeRF TensoRF supasses the original TensoRF, indicating that Mesh2NeRF supervision contributes to the neural network’s ability to learn accurate scene representation. We present qualitative results from the Sketchfab dataset in Fig.11. Renderings from two views for each object are compared with NeRF baselines, accompanied by corresponding PSNR values. Similarly, Fig.12 showcases qualitative results on the ABO dataset, with renderings from two views for each object compared alongside corresponding LPIPS values. These figures demonstrate that Mesh2NeRF NeRF outperforms NeRF baselines.